# Absence of censoring inequalities in random quantum circuits

### v1

@@ V Conclusion
We give several examples of architecture which violate censoring inequalities. However, all of our architectures feature a connectivity graph with nontrivial topology. It thus remains plausible that censoring inequalities do hold for architectures with geometric locality on treelike graphs, since on these graphs it is more difficult for information to “dodge” scrambling. Our examples also feature discrete unitary gates. It may be interesting to investigate a continuous-time version of this problem, such as the conditions under which adding a new interaction to a random Hamiltonian can slow down thermalization.

### v2

@@ V Conclusion
Our results suggest a more general question about how global scrambling depends on local dynamics. In the approximate $t$-design setting, Ref. 15 gives a relationship between local and global spectral gaps, suggesting that the rate of global scrambling is at worst proportional to the rate of local scrambling. In the same vein, Ref. 16 shows that certain highly-entangling local ensembles give $t$-designs faster than Haar-random local gates. One can interpret this result as showing, loosely, that removing randomness from the circuit in a way that increases the typical amount of local scrambling makes the circuit a better approximate $t$-design. In this framing, our results show that even removing randomness in a way that decreases local scrambling can have the same effect. It remains unclear if there is any clear conceptual description of the relationship between local and global behavior.
@@
We give several examples of architecture which violate censoring inequalities. However, all of our architectures feature a connectivity graph with nontrivial topology. However, none of our architectures feature a linear connectivity graph. It thus remains plausible that censoring inequalities do hold for architectures with geometric locality on the line, since in this case it is more difficult for information to “dodge” scrambling. Our examples also feature discrete unitary gates. It may be interesting to investigate a continuous-time version of this problem, such as the conditions under which adding a new interaction to a random Hamiltonian can slow down thermalization.